# ELITR-Bench: A Meeting Assistant Benchmark for Long-Context Language Models

### v2

@@ B.2 Additional results on question type and answer position
In this section, we provide the full results per question type and answer position to expand the compact results of the GPT and LLaMA-2 model families given in Section 5.2. The results are given in Table 9 and were obtained on ELITR-Bench-QA’s test set in the single-turn setting. Looking at the global model performance over the different question types and answer positions, we do not identify any clear trend highlighting a question type or position answer as notably easier or harder.
In constrast, past work reported a “lost in the middle” effect [29], stating that the middle of a model’s context tends to be overlooked more often than the beginning or end of the context. To further investigate this phenomenon in our dataset, we conducted a statistical hypothesis test on the scores obtained by each individual model. Specifically, we ran a one-tailed Welch’s t-test [43] with the following alternative hypothesis: “The average score for questions with middle-position answers is lower than the average score of other questions”. The p-values obtained for each model’s set of scores are given in Table 10. Interestingly, we observe that the “lost in the middle” hypothesis is statistically verified (p-value <<< 0.05) for only two models: LongChat-7B-v1.5 (p-value === 0.032) and Vicuna-7B-v1.5 (p-value === 0.046). While we do not have a clear explanation about which of these two models’ characteristics caused that effect, these models have in common that they are based on LLaMA-2-7B and were trained by the same LMSYS organization [11, 25]. It is then possible – although purely hypothetical – that the specific fine-tuning recipe followed by LMSYS on LLaMA-2-7B for these two models led to the “lost in the middle” effect.
B.3 Updated results with recent long-context LLMs
In this section, we provide updated results to include some of the latest long-context LLMs that have been released after the core experiments of this paper have been conducted. In particular, we added the three following models: OpenAI’s GPT-4o, [^17] https://platform.openai.com/docs/models/gpt-4o GradientAI’s LLaMA-3-8B-Instruct-262K [^18] https://huggingface.co/gradientai/Llama-3-8B-Instruct-262k which was fine-tuned from the original LLaMA-3 model, [^19] https://llama.meta.com/llama3/ and Microsoft’s Phi-3-Small-128K-Instruct. [^20] https://huggingface.co/microsoft/Phi-3-small-128k-instruct For GPT-4o, we used the same hyperparameters as for GPT-4 and GPT-4.5: nucleus sampling with a temperature of 0.6 and top-p of 0.9. For LLaMA-3-8B-Instruct-262K and Phi-3-Small-128K-Instruct, inference was done by greedy decoding and we used their original chat templates. QA markers and repetition penalty have not been included.
The comparison of these additional models and previously used approaches is provided in Table 11. These results show that GPT-4o performed essentially on par with GPT-4, with a small gain on the dev set. Interestingly, LLaMA-3-8B-Instruct-262K remains close in performance to LLaMA-2 models and only slightly outperforms the best among these, while Phi-3-Small-128K-Instruct beats both LLaMA-2 and LLaMA-3 models by a good margin. It even outperforms GPT-3.5 despite having only 7B parameters. This suggests that this Phi-3 model is a good open-source alternative to more powerful models such as GPT-4 and GPT-4o for our meeting assistant task.
Table 11: Results on single-turn ELITR-Bench-QA updated with recent long-context LLMs: GPT-4o, LLaMA-3-8B-Instruct-262K, and Phi-3-Small-128K-Instruct. The reported numbers correspond to the average scores from 1 to 10 (higher is better) obtained by a GPT-4 evaluator, on a single seeded run for the dev set and 3 seeded runs for the test set. Boldface numbers correspond to the best performance among proprietary or open-source models.
Model Single-turn ELITR-Bench-QA
Dev Test Mean
GPT-3.5 7.04 7.44 7.24
GPT-4 8.21 8.39 8.30
GPT-4o 8.50 8.44 8.47
LongAlpaca-7B 5.89 5.60 5.75
LongAlpaca-13B 6.17 6.25 6.21
LongChat-7B-v1.5 6.60 5.78 6.19
Vicuna-7B-v1.5 5.42 5.61 5.51
Vicuna-13B-v1.5 5.92 6.52 6.22
LongAlign-7B 6.11 6.46 6.28
LongAlign-13B 6.27 6.33 6.30
LLaMA-3-8B-Instruct-262K 6.83 6.51 6.67
Phi-3-Small-128K-Instruct 7.37 7.34 7.36
Appendix C Crowdsourcing study details
@@
The guidelines provided for this study start with general information about the task as well as the 10-point score rubric given in Fig. 6, in order to help annotators calibrate their scores with concrete criteria. Then the interface presents a tuple composed of a question, its ground-truth answer, and an LLM response to evaluate. From this tuple, the annotator is asked to grade the LLM response with a score ranging from 1 to 10, following the provided score rubric. A screenshot of our interface is shown in Fig. 3.
@@
Appendix D ELITR-Bench excerpt
@@
We provide in Table 13 an excerpt of meeting 010 from the dev set of the ELITR corpus [33]. Entities, such as (PERSON10), (PERSON19), and [ORGANIZATION11], have been de-identified in the original work for the sake of anonymization. Below the excerpt, we provide 4 questions (and their respective answers) related to the same meeting, which have been added through the proposed ELITR-Bench. For each question, we indicate its type between brackets (i.e., Who, What, When, or How many).
Appendix E Prompts
@@
In this section, we list the different prompts used in the paper, both for response generation and evaluation. The prompt for response generation follows the same general template given in Fig. 4 for every evaluated model – both proprietary and open-source models. Then, questions and answers are appended to the prompt as described in Section 4 – either a single question per conversation in the single-turn mode, or all the questions of a meeting in sequence in the multi-turn mode. As detailed in Appendix A.2, we slightly modify this base prompt depending on the model-specific selected configuration. As a reminder, these alterations may take two forms: the use of a chat template (which only adds special tags to the prompt) and the use of question-answer markers (which add ‘QUESTION:’ before a question and ‘ANSWER:’ before an answer).

### v3

@@ C.3 Investigation of a “lost in the middle” effect
LLaMA-3.1-8B 0.308
Phi-3-small 0.541
Past work reported a “lost in the middle” effect [31], stating that the middle of a model’s context tends to be overlooked more often than the beginning or end of the context. To further investigate this phenomenon in our dataset, we conducted a statistical hypothesis test on the scores obtained by each individual model. Specifically, we ran a one-tailed Welch’s t-test [46] with the following alternative hypothesis: “The average score for questions with middle-position answers is lower than the average score of other questions.” The p-values obtained for each model’s set of scores are given in Table 10. Interestingly, we observe that the “lost in the middle” hypothesis is statistically verified (p-value <<< 0.05) for only two models: LongChat-7B-v1.5 (p-value === 0.032) and Vicuna-7B-v1.5 (p-value === 0.046). While we do not have a clear explanation about which of these two models’ characteristics caused that effect, these models have in common that they are based on LLaMA-2-7B and were trained by the same LMSYS organization. It is then possible – although purely hypothetical – that the specific fine-tuning recipe followed by LMSYS on LLaMA-2-7B for these two models led to the “lost in the middle” effect.
Appendix D Crowdsourcing study details
@@
The guidelines provided for this study start with general information about the task as well as the 10-point score rubric given in Fig. 7, in order to help annotators calibrate their scores with concrete criteria. Then the interface presents a tuple composed of a question, its ground-truth answer, and an LLM response to evaluate. Given this tuple, the annotator is asked to grade the LLM response with a score ranging from 1 to 10, following the provided score rubric. A screenshot of our interface is shown in Fig. 4.
@@
Appendix E ELITR-Bench excerpt
@@
We provide in Table 12 an excerpt of meeting 010 from the dev set of the original ELITR corpus [35]. Entities, such as (PERSON10), (PERSON19), and [ORGANIZATION11], have been de-identified in the original work for the sake of anonymization. Below the excerpt, we provide 4 questions (and their respective answers) related to the same meeting, which have been added through the proposed ELITR-Bench. For each question, we indicate its type between brackets (i.e., Who, What, When, or How many).
Appendix F Prompts
@@
In this section, we list the different prompts used in the paper, both for response generation and evaluation. The prompt for response generation follows the same general template given in Fig. 5 for every evaluated model – both proprietary and open models. Then, questions and answers are appended to the prompt as described in Section 4 – either a single question per conversation in the single-turn mode, or all the questions of a meeting in sequence in the multi-turn mode. As detailed in Appendix B.2, we slightly modify this base prompt depending on the model-specific selected configuration. As a reminder, these alterations may take two forms: the use of a chat template (which only adds special tags to the prompt) and the use of question-answer markers (which add ‘QUESTION:’ before a question and ‘ANSWER:’ before an answer).